• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name brads

    Log entry time 17:44:01 on February 26, 2009

    Entry number 263530

    This entry is a followup to: 263511

    keyword=BB Cerenkov gains checked -- RUNLIST

    Most gains looked reasonable. PMTs 2, 12 needed a significant change. Data are in the files listed below.

    PMT 17 is pretty noisy. (Been true for a while. We didn't want to swap it out since it will require a longish down to flush the entire Cerenkov tank.)

    HVs for PMTs 2*,4,8,9,10,12* were tweaked.
    
    (*) changed by >50V
    
    Previous HVs recorded here (run 1810):
      http://www.jlab.org/~adaq/halog/html/0902_archive/090225223803.html
    
    New HVs can be found here (run 1841):
      http://www.jlab.org/~adaq/halog/html/0902_archive/090226170318.html
    
    Run 1829:
      - HVs for 1--10 on, pulser on
      - LED #11, #15, #20 plugged in
        - LED drive pulse ampl. set to provide 1--2 p.e. pulse on the scope
    for PMT 10
      - Pedestal run
        - T8 trigger dominated (others prescaled away)
      = peds:  483, 407, 398, 479, 543, 499, 603, 527, 460, 465
    
    Run 1830:
      - HVs for 1--10 on, pulser on
      - LED #11, #15, #20 plugged in
      = aligning 1--8  to pedestal + 30
      = aligning 9--10 to pedestal + 60 (since they are not split)
    
      = adjusted PMTs 2,3,4,8
      = adjusted PMTs 9, 10
    
    Run 1831:
      - HVs for 1--10 on, pulser on
      - LED #11, #15, #20 plugged in
      = adjusted PMT 4
    
    Run 1832:
      - HVs for 1--10 on, pulser on
      - LED #11, #15, #20 plugged in
      = adjusted PMT 4
    
    Run 1833:
      - HVs for 1--10 on, pulser on
      - LED #11, #15, #20 plugged in
        - raised LED drive voltage a bit to enhance 2 p.e. peaks
      = looks OK
    
    
    Run 1834:
      - HVs for 11--20 on, pulser on
      - LED #1, #5, #10 plugged in
        - adjusted LED drive voltage for this LED group
      - Pedestal run
        - T8 trigger dominated (others prescaled away)
      = peds: 303, 425, 515, 578, 579, 443, 410, 542, 329, 412
    
    Run 1835:
      - HVs for 11--20 on, pulser on
      - LED #1, #5, #10 plugged in
      = adjusted 12
    
    Run 1836:
      - HVs for 11--20 on, pulser on
      - LED #1, #5, #10 plugged in
        - LED drive was too large, lowered
      = 12 good
    
    Run 1837:
      - HVs for 11--20 on, pulser on
      - LED #1, #5, #10 plugged in
        - LED drive was still too large for some of the PMTs, lowered it again
      = Gains look pretty good.
      = PMT 17,18 have much noisier pedestals than the rest, will investigate
    that
        now.